OpenStudy (anonymous):

Find the coordinates of the x and y intercepts for this equation. y = 2 x - 3 x -intercept (____ , 0) y -intercept (0,____ ) Find the coordinates of the x and y intercepts for this equation. y = 2 x - 3 x -intercept (____ , 0) y -intercept (0,____ ) @Mathematics

OpenStudy (anonymous):

solve for x intercept 0=2x-3 3=2x x=1.5 solve for y intercept y=2(0)-3 y=-3

OpenStudy (anonymous):

Ok so the answer is X intercept = 1.5 Y intercept = -3